OpenStudy (anonymous):

the flower garden in bay area park measures 24 meters by 30 meters. the garden is surrounded by a paved walk of uniform width. if the combined area of the paved walk and the garden is 1080m^2, what is the width of the walk?

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

24*30 + 2*x*24 + 2*x*30 - 4x^2 = 1082 You have to solve: 108*x - 4*x^2 = 1082-720 = 362

OpenStudy (anonymous):

AKA 4x^2-108x+362=0

OpenStudy (anonymous):

Furthermore: 4x^2-108x+362=0 <=> (2x-27)^2=367 <=> 2x-27=±√367 <=> 2x=27±√367 <=> x=(27±√367)/2

OpenStudy (anonymous):

Approximately: x=23.08 AND x=3.92 The width of the walkway is not the width of one of the sides, therefore the answer is the second one. The paved walk is 3.92 meters wide.